I'm having trouble writing equations that models how old in years each of us will be, when our ages add up to 150 years old. I had to use three people's ages 8,28,30. Explain the reasoning which helped you develop your equation?
Answer by stanbon(75887)

I'm having trouble writing equations that models how old in years each of us will be, when our ages add up to 150 years old. I had to use three people's ages 8,28,30. Explain the reasoning which helped you develop your equation?
----
Those three person's ages do not add up to 150 years old.
---
You might use 40,60, and 50 but not 8,28 and 30.
----
If you use a = 40, b = 60, and c = 50
you would get 3 equations:
---
a + b + c = 150
a = b-20
c = b+10
---------
That's an example.
----------------------------
Cheers,
Stan H.
====================